# SynLogic: Synthesizing Verifiable Reasoning Data at Scale for Learning Logical Reasoning and Beyond

### v2

@@ 2.2 The Data Synthesis Framework
Risk of Data Contamination
@@ 3.2 Results
Substantial Improvements in Logical Reasoning
@@
Generalization to Mathematical Domains
@@
Increased Chain-of-Thought Length
@@ 4.1 Setup Details
For mathematical training data, we directly utilize the 17k samples provided in DAPO (Yu et al., 2025). For coding data, we assembled approximately 9k samples from various online coding platforms such as Codeforces. We adapted a similar prompt template to that used in our SynLogic training; the detailed templates are presented in Appendix B.1. We maintained the same reward design approach as described in §3.1. Specifically, for coding tasks, the reward is set as 1 if the output is correctly formatted and all test cases pass; otherwise, the reward is 0.
@@ 4.2 Mixing SynLogic with Math or Code Data: A Pilot Ablation Study
Mixed Training with Math
@@
Mixed Training with Code
@@ 5 Conclusion
Limitations
First, due to limited computational resources, we do not conduct precise difficulty tuning for every task, meaning the curated data may not be fully optimized for training 7B and 32B models. Second, we do not implement training with dynamically adjustable difficulty, where sample complexity progressively increases during training. We hypothesize that such an approach could further enhance the logical reasoning capabilities of LLMs. Both of these directions are left as opportunities for future exploration.
@@ A.1 Task Composition and Sources
Table 4 presents the diverse collection of tasks incorporated in SynLogic. We have carefully selected these tasks from established benchmarks including KOR-Bench [Ma et al., 2024], BBH [Suzgun et al., 2022], and BBEH [Kazemi et al., 2025]. Additionally, we integrated some logical reasoning tasks not previously featured in these benchmarks, such as Mathador [Kurtic et al., 2024] and Minesweeper [Li et al., 2024].
Our collection comprises 35 distinct tasks, with only two (Zebra Puzzle [Lin et al., 2024] and ARC-AGI [Chollet, 2019]) using existing data sources. For all remaining tasks, we generated custom datasets. Importantly, we developed and implemented verifiers for all tasks in the collection, ensuring consistent evaluation across the benchmark.

### v4

@@ 2.2 The Data Synthesis Framework
2.2.1 Risk of Data Contamination
@@ 3.2 Results
3.2.1 Substantial Improvements in Logical Reasoning
@@
3.2.2 Generalization to Mathematical Domains
@@
3.2.3 Increased Chain-of-Thought Length
@@ 4.1 Setup Details
For mathematical training data, we directly utilize the 17k samples provided in DAPO (Yu et al., 2025). For coding data, we assembled approximately 9k samples from various online coding platforms such as Codeforces. We adapted a similar prompt template to that used in our SynLogic training; the detailed templates are presented in Appendix B.1. We maintained the same reward design approach as described in §3.1.2. Specifically, for coding tasks, the reward is set as 1 if the output is correctly formatted and all test cases pass; otherwise, the reward is 0.
@@ 4.2 Mixing SynLogic with Math or Code Data: A Pilot Ablation Study
4.2.1 Mixed Training with Math
@@
4.2.2 Mixed Training with Code
@@ A.1 Task Composition and Sources
Table 4 presents the diverse collection of tasks incorporated in SynLogic. We have carefully selected these tasks from established benchmarks including KOR-Bench (Ma et al., 2024), BBH (Suzgun et al., 2022), and BBEH (Kazemi et al., 2025). Additionally, we integrated some logical reasoning tasks not previously featured in these benchmarks, such as Mathador (Kurtic et al., 2024) and Minesweeper (Li et al., 2024).
Our collection comprises 35 distinct tasks, with only two (Zebra Puzzle (Lin et al., 2024) and ARC-AGI (Chollet, 2019)) using existing data sources. For all remaining tasks, we generated custom datasets. Importantly, we developed and implemented verifiers for all tasks in the collection, ensuring consistent evaluation across the benchmark.